OpenStudy (anonymous):

There is a sphere that has a radius of 9 inches. The sphere has a smaller sphere inside that has a radius of 3 inches. What is the ratio of the volume of the small sphere to the volume of the larger sphere?

jimthompson5910 (jim_thompson5910):

What is the volume of the larger sphere?

OpenStudy (anonymous):

you dont know it

jimthompson5910 (jim_thompson5910):

Are you able to find it?

OpenStudy (anonymous):

I dont know how

jimthompson5910 (jim_thompson5910):

You would use the formula \[\Large V = \frac{4}{3}\pi*r^3\]

jimthompson5910 (jim_thompson5910):

for the larger sphere, r = 9 the smaller one, r = 3

OpenStudy (anonymous):

So it would be V= 4/3 pie (729) how would you multiply 4/3

jimthompson5910 (jim_thompson5910):

think of 729 as 729/1

jimthompson5910 (jim_thompson5910):

so (4/3)*729 = (4/3)*(729/1) = (4*729)/(3*1) = 2916/3 = 972

OpenStudy (anonymous):

ohh! ok thank you!

jimthompson5910 (jim_thompson5910):

meaning that \[\Large V = \frac{4}{3}\pi*(729)\] becomes \[\Large V = 972\pi\]

jimthompson5910 (jim_thompson5910):

do the same for the smaller sphere, then you'll have your ratio (but don't forget to fully reduce this ratio)

OpenStudy (anonymous):

what would the ratio be to find the surface area of the two spheres?

jimthompson5910 (jim_thompson5910):

what did you get for the ratio of the volumes?

OpenStudy (anonymous):

27 over 1, and the pie cancels out

jimthompson5910 (jim_thompson5910):

they want it in the form (small volume):(large volume) and you reduce that

jimthompson5910 (jim_thompson5910):

so that ratio would be 1:27 when fully reduced

jimthompson5910 (jim_thompson5910):

as for the surface area, you use the formula \[\Large SA = 4\pi*r^2\]

OpenStudy (anonymous):

oh yes I switched the rations, and ok thanks i will use that for SA